# Coordination-free Collaborative Replication based on Operational Transformation

### v1

@@ 2 Collaborative Replication
Although the research on CRDTs is vast and multifaceted, CRDTs and CCR can be seen within the same framework in that they are updated within sites and reflected as shared data. However, CCR has more concise and useful aspects than the basic structure of CRDTs, and this is the focus of the discussion in this paper.
@@ 2.0.1 Shattering the CALM for Collaborative Data Sharing
However, monotonicity is not the only golden rule for collaborative data sharing in distributed systems! A binary operation is commutative if the order of its operands makes no difference to the result. The confluence property as applied to program operations is a generalization of commutativity.
@@ 2.0.2 Operational Transformation is Disliked in Replication, but …
Our CCR is a new challenge to exploit OT-based data sharing. It comes from the Tutorial Blog [^1] http://blog.haskell-exists.com/yuras/posts/realtime-collaborative-editor.html “Realtime collaborative editor. Algebraic properties of the problem” by Yuras Shumovich, March 5, 2016 which explain concisely the idea and pitfalls of Collaborative Editor (RTCE) implementations based on operational transformation. Although it assumes that the system uses a server for maintaining the state of clients, it works in a server-less environment, too.
@@ 3.1.2 Operational Transformation
[…]
@@
  • The Compositional Property is very natural for the transformation is defined for the basic operations and extends it to accept any operations of the monoid $(O,!,\odot)$ any element
@@ 4.0.4 Eventual Consistency of Collaborative Replication
The above observation shows that the replication process should terminate when OT produces the identity operation.

### v2

@@ 2.0.1 Shattering the CALM for Collaborative Data Sharing
However, monotonicity is not the only golden rule for coordination-free data sharing in distributed systems! A binary operation is commutative if the order of its operands makes no difference to the result. The confluence property as applied to program operations is a generalization of commutativity.
@@ 2.0.2 Operational Transformation is Disliked in Replication, but …
Our CCR is a new challenge to exploit OT-based data sharing. It comes from the Tutorial Blog [^1] http://blog.haskell-exists.com/yuras/posts/realtime-collaborative-editor.html “Realtime collaborative editor. Algebraic properties of the problem” by Yuras Shumovich, March 5, 2016 which explains concisely the idea and pitfalls of Collaborative Editor (RTCE) implementations based on operational transformation. Although it assumes that the system uses a server for maintaining the state of clients, it works in a server-less environment, too.
@@ 3.1.2 Operational Transformation
Local Confluence Property The transformation $T_{D}$ takes operations $p$ and $q$ on $D$ and produces $(p^{\prime},q^{\prime})=T_{D}(p,q)$ which makes $D\odot p\odot q^{\prime}=D\odot q\odot p^{\prime}$. That is, $p$ and $q$ on the same $D$ are made into the common confluent state by $T_{D}$. This property is called “TP1-Confluence” which is the basis of “TP2-Confluence” for global integration stability described below.
@@
  • The Compositional Property is very natural for the transformation is defined for the basic operations and extends it to accept any operations of the monoid $(O,!,\odot)$.